OpenStudy (anonymous):

use the properties of exponents to simplify. write with positive exponents. Assume that all variables represent positive real numbers. use the properties of exponents to simplify. write with positive exponents. Assume that all variables represent positive real numbers. @Mathematics

OpenStudy (anonymous):

please show how?

OpenStudy (anonymous):

\[b^ab^a = b^{a+a}\]

OpenStudy (anonymous):

\[\frac{b^a}{b^c}=b^{a-c}\]

OpenStudy (anonymous):

-b